# Masked Sensory-Temporal Attention for Sensor Generalization in Quadruped Locomotion

### v1

@@ Abstract
With the rising focus on quadrupeds, a generalized policy capable of handling different robot models and sensory inputs will be highly beneficial. Although several methods have been proposed to address different morphologies, it remains a challenge for learning-based policies to manage various combinations of proprioceptive information. This paper presents Masked Sensory-Temporal Attention (MSTA), a novel transformer-based model with masking for quadruped locomotion. It employs direct sensor-level attention to enhance sensory-temporal understanding and handle different combinations of sensor data, serving as a foundation for incorporating unseen information. This model can effectively understand its states even with a large portion of missing information, and is flexible enough to be deployed on a physical system despite the long input sequence.
11footnotetext: NVIDIA AI Technology Centre (NVAITC); e-mail: {dikail,jianxiongy,ssee}@nvidia.com22footnotetext: College of Computing and Data Science, Nanyang Technological University, Singapore; e-mail: dikai001@e.ntu.edu.sg, tianwei.zhang@ntu.edu.sg33footnotetext: also with Nanyang Technological University and Coventry University
Keywords: Quadruped robots, Locomotion, Transformer
1 Introduction
@@
Quadrupedal robots have showcased their capability to navigate in various challenging terrains, thanks to the rapid advancements of deep reinforcement learning (RL) technology [1, 2, 3, 4]. With the increasing availability of affordable quadruped robots on the market, there is a growing interest in developing a general-purpose locomotion policy that can fit all types of quadrupedal devices. Typically learned locomotion policies are specific to the robot, observation space, and task they were trained on, making it challenging to transfer or generalize to other robots or scenarios.
@@
Recent works, such as GenLoco [5] and ManyQuadrupeds [6], have successfully developed a single policy with the ability to adapt to diverse morphologies and degrees of freedom (DoF). However, these models depend on a fixed observation space input to generate latent space representations. They become ineffective when facing the following situations: (1) deployed to quadrupeds equipped with a different sensor set, (2) sensor data turn to unreliable due to wear and tear, (3) new input is needed for new task. Since sensory feedback is interrelated and each sensor plays a critical role at different stages of locomotion [7], a policy with a deep understanding of proprioceptive information to handle flexible input is desired, to enhance the generalization, flexibility, and extensibility.
@@
One promising solution is self-attention-based transformers [8], which have demonstrated exceptional capabilities in understanding complex sequential information of arbitrary lengths. They have been widely used in robotics to enhance various tasks with multimodal processes [9, 10, 11, 12]. However, due to the complex model structures and vast parameters, robots driven by transformers often run at very low frequency [10, 11], or depend on external high-power computing platforms [13]. To address these issues, two common strategies are proposed to achieve smooth and stable control for transfermer-based locomotion tasks. The first strategy involves outputting high-level commands as an interface, such as gait pattern [14] and velocity commands [15]. An external joint controller, which requires additional design and training, is then used to convert these high-level commands into joint commands. However, the application of this method is restricted to the policy combined with perception and human-robot interaction. The second strategy, targeting the vanilla locomotion tasks, applies linear projections to observation-action data for tokenization [16, 17, 18]. Although this method is straightforward and efficient for producing joint commands in an end-to-end manner, it limits the transformer’s direct access to sensory information and still relies on fixed input, thereby constraining its in-context understanding capability and multimodal nature of the data.
@@
To address the above limitations, we propose Masked Sensory-Temporal Attention, a novel transformer-based model for end-to-end quadruped locomotion control. It achieves sensor input generalization with its multimodal nature, while still being directly deployable on physical systems. Specifically in Masked Sensory-Temporal Attention, all sensory data are discretized and tokenized to form a long proprioceptive information sequence. Inspired by the work [19] on learning spatiotemporal information in video understanding, a random mask is applied to remove a portion of the observation during training. This significantly enhances the model’s sensory-temporal understanding, to better handle different combinations of sensor data and serve as a foundation for incorporating unseen data. Additionally, it aids in identifying the most essential sensory information, thereby reducing the computational power required for physical deployment. Extensive experiments conducted in simulation demonstrate that our method can efficiently handle incomplete sensory information, even with half of the data missing. It is also robust against unseen data, making it a solid foundation for further extensions. With direct sensory-temporal attention, the model is flexible enough to mix-and-match desired information for finetuneing, meeting the requirement for different end-to-end quadruped locomotion control in the physical world.
2 Related Work
2.1 Reinforcement Learning in Legged Locomotion
@@
Reinforcement learning (RL) has gained significant attention in developing robotic controllers for tasks such as legged locomotion [1, 2, 3, 4, 16], eliminating the need for extensive prior knowledge. With the advance of robotic simulation, RL-based locomotion is often trained in virtual environments [4, 20, 21] with diverse terrains [21] and randomized environmental factors [3, 4] to improve the robustness of the agent. This technique is commonly known as domain randomization (DR). Training in simulators also provides rich information, some of which is not easily accessible in the real world (i.e., privileged information). To better interpret such information and bridge the sim-to-real gap when deploying policies to the physical world, system identification is commonly used to transfer knowledge to a deployable student policy. Lee et al. [3] employed action imitation to infer teacher behaviors using historical proprioceptive data. Kumar et al. [4] further developed a two-stage adaption framework for faster and more robust online transfer, which has become the foundation for many subsequent works [22, 16]. Another approach combines transfer loss with RL loss for joint optimization [23, 17] to allow the student to explore freely to maximize the reward return.
2.2 Transformer in Robotics
@@
New models and methods are proposed to leverage the power of transformers in RL and robotics. For instance, Decision Transformer [24] converts states, actions, and rewards into embeddings using an encoder. Trajectory Transformer [25] uses the complete discretized trajectory for language model-like autoregressive prediction. Building on these frameworks, Gato [9] was developed to serve as a generalist agent for hundreds of tasks, including real-world robotic manipulation. Vision Language Models (VLM) [10] and Vision Language Action Models (VLA) [11, 12] use transformers as interfaces for scene and language understanding to provide high-level commands for robotic control and human-robot interaction.
@@
For legged locomotion, Yang et al. [15] developed a transformer model for vision-based locomotion. It outputs high-level velocity commands and relies on a dedicated low-level controller for motor control. To bring transformers to direct motor control, Lai et al. [16] proposed TERT, which utilizes historical observation-action pairs to generate target motor commands directly. Radosavovic et al. [17] applied a similar method to the bipedal locomotion task and later reformulated it as a next token prediction problem [18].
3 Preliminary
@@
We adopt the two-stage teacher-student transfer approach from TERT [16] as the basis of our framework. It utilizes a well-trained teacher policy through RL with privileged information.
3.1 Simulation Environment
@@
We implement the simulation environment based on Isaac Gym and its open source library IsaacGymEnvs [20] to enable massive parallel training.
Terrain and Curriculum. We adopt the terrain curriculum from [21] to achieve locomotion in dynamic terrains. Five terrain types (smooth slope, rough slope, stairs up, stairs down, discrete obstacle) are used to spawn a total of 20 terrains with proportions of $[0.1,0.1,0.35,0.25,0.2]$ respectively. Each terrain has 10 levels with the increasing difficulty for larger slope angle, higher stairs step and larger obstacle size. During training, the agent progresses to the next level when the tracked linear reward reaches 80% of the maximum achievable value and regresses if it fails to reach 25%. If any agent finishes the highest level, it will be sent to a random level to continue exploring.
Domain Randomization. To enhance the robustness of the policy, DR is used in the simulation. The commanded longitudinal and lateral velocity is sampled in [-1.0, 1.0] m/s and horizontal angular velocity in [-1.0, 1.0] rad/s. The command is resampled every 5 seconds without any curriculum. Due to the significant computational power and time required for transformer calculations, system delay is introduced to simulate the delay in action updates. Please refer to the Appendix for the complete DR parameters.
@@
Observations and Actions. The privilege observation $e_{t}$ for teacher training contains ground truth data gathered fro simulation, including base linear and angular velocity, orientation, surrounding height map and randomized parameters as described above. For proprioceptive information, we use three commonly seen low-level sensors from quadrupeds, joint encoders, IMU and foot contact sensors. These sensors can provide five sensory data, including joint position $q\in\mathbb{R}^{12}$, joint velocity $\dot{q}\in\mathbb{R}^{12}$, angular velocity $\omega\in\mathbb{R}^{3}$, gravity vector $g\in\mathbb{R}^{3}$ and binary foot contact $c\in\mathbb{R}^{4}$. Furthermore, the randomly sampled user command target $cmd=[v_{x},v_{y},\omega_{z}]$ and actions from previous step $a_{t-1}\in\mathbb{R}^{12}$ is added, resulting in a observation of $o_{t}\in\mathbb{R}^{49}$ for each step. To gather the temporal information, a list of historical proprioceptive information $[o_{t},o_{t-1},\cdots,o_{t-T}]$ from past $T=15$ steps is stored and passed to the student network. Thus, full observation is in the $\mathbb{R}^{49\times 15}$ space. Both teacher and student outputs the desired joint position $a_{t}$, which is further processed by a PD controller for the output torque $\tau=K_{p}(\hat{q}-q)+K_{d}(\hat{\dot{q}}-\dot{q})$, with base stiffness and damping set to 30 and 0.7 respectively and the target joint velocity $\hat{\dot{q}}$ set to 0.
@@
Reward Function for RL. The reward functions are designed to encourage the agent to follow the commanded velocity. Following [21, 4, 26], we primarily penalize linear and angular movement along other axes, large joint acceleration and excessive power consumption. Please refer to the Appendix for the complete reward structure.
3.2 Teacher Policy and Training
@@ 4 Methodology
We present MSTA, a novel transformer-based framework to generate a generalized understanding of low-level proprioceptive information for quadruped locomotion in complex terrains. With this foundational understanding, this framework is capable of handling different combinations of sensory inputs, enabling better generalization and flexibility. It can potentially be extended to incorporate high-dimensional sensors for more complex tasks. Figure 1 shows the overview of MSTA.
@@
Sensory Data Tokenize. To avoid using linear projection like previous transformer-based locomotion controllers [16, 18], we need to map continuous data to tokens directly. Following previous works [9, 10, 11], we first normalize the sensory data based on the collected mean and variance, and then pass them through an encoder. This is similar to Gato [9], which discretizes the value into 256 bins. The data are further mapped into a learnable embedding space with 128 dimensions.
@@
Positional Embedding and Sensor Type Embedding. To distinguish proprioceptive information from different sensors with multiple dimensions and historical relations, two additional embeddings are added. The first one is a fixed 2D sin-cos position embedding [28] applied on the sensor dimension and time axis. This allows the framework to handle sensors with varying lengths of dimensions and historical time windows and be easily extendable. To accommodate the multimodal nature of the sensory data, another learnable embedding is add to indicate each sensor type. This enables easy mix and match of information from different sensors without concerns about the order or placeholders.
Figure 1: Overview of our framework. We gather proprioceptive information from commonly seen low-level sensors for discretization and tokenization. Similar to video understanding, we add additional embedding in three dimensions: sensor type, sensor dim and time. Before being passed to the transformer, a random mask is applied to partially remove the information and a learnable state embedding <S> is used to consolidate the information for action prediction. The target joint position output is passed to the PD controller for direct joint control.
@@
Random Masking. Inspired by the use of masking in image and video understanding [29, 19], we randomly mask out portions of the sensory information. This brings two benefits. First, since only part of the sensory data are visible to the network, the model is required to infer and reconstruct the missing information from them, thereby enhancing its understanding of the relationships between different sensory inputs. Second, random masking significantly reduces the training time and computational resources required, as the complexity of self attention is necessarily quadratic in the input length, making it more feasible to run in massive parallel.
@@
Transformer Model. We implement a transformer model to process the generated tokens. The model consists of multi-head self-attention blocks with a MLP ratio of 2.0. An additional learnable state embedding <S> is added to consolidate the processed information [30], which is subsequently projected into the action space with a MLP layer $\pi$:
$\begin{gathered}l_{t}=\mathrm{MSTA{}}([o_{t},o_{t-1},\cdots,o_{t-T}]),\;\;\;\;a_{t}=\pi(l_{t})\end{gathered}$ (2)
@@
Teacher-Student Transfer. Following TERT [16], we train MSTA with a two-stage transfer strategy. In the first offline pretraining stage, trajectory is gathered by unrolling the well-trained teacher policy while the student will predict the next actions. This is to ensure that the student can produce reasonable actions during the second online correction stage to overcome the gap of distribution shift by training on its own trajectory. In each stage, 50 million trajectories are collected to train the transformer for 80k updates. We minimize the loss for the next step prediction:

### v2

@@ Abstract
With the rising focus on quadrupeds, a generalized policy capable of handling different robot models and sensor inputs becomes highly beneficial. Although several methods have been proposed to address different morphologies, it remains a challenge for learning-based policies to manage various combinations of proprioceptive information. This paper presents Masked Sensory-Temporal Attention (MSTA), a novel transformer-based mechanism with masking for quadruped locomotion. It employs direct sensor-level attention to enhance the sensory-temporal understanding and handle different combinations of sensor data, serving as a foundation for incorporating unseen information. MSTA can effectively understand its states even with a large portion of missing information, and is flexible enough to be deployed on physical systems despite the long input sequence.
I Introduction
@@
Benefiting from the rapid advancements of deep reinforcement learning (RL) technology [1, 2, 3, 4], quadrupedal robots have showcased their capability to navigate in diverse complex terrains. With the increasing availability of affordable quadruped robots on the market, there is a growing interest in developing general-purpose locomotion policies that can fit all types of quadrupedal devices. Unfortunately, existing learning-based locomotion policies are trained for specific models, observation spaces, and tasks, making it challenging to transfer or generalize to other unseen robots or scenarios.
@@
Recently, researchers have developed some generalized policies for quadruped locomotion, such as GenLoco [5] and ManyQuadrupeds [6], which have the ability to adapt to diverse morphologies. However, these methods still depend on a fixed observation space input for generating latent space representations. They become ineffective when facing the following situations: (1) deployment on quadrupeds with a different sensor set; (2) unreliable sensor data due to wear and tear, (3) adapting to a new task with new input. Since sensory feedback is interrelated and each sensor plays a critical role at different stages of the locomotion [7], a policy with a deep understanding of proprioceptive information to handle flexible inputs is desired, to enhance the generalization, flexibility, and extensibility.
Figure 1: Commonly seen low-level sensors on a quadrupedal robot. However, actual sensor set is still different across models, and sensor degradation can cause part of sensor data to be unreliable or even unavailable. With MSTA, we create a generalized model to enhance the understanding of sensor information to handle variable sensor input for quadruped locomotion.
@@
One promising solution is self-attention-based transformers [8], which have demonstrated exceptional capabilities in understanding complex sequential information of arbitrary lengths. They have been widely used in robotics to enhance various tasks with multimodal processes [9, 10, 11, 12]. However, due to the complex model structures and vast parameters, robots driven by transformers often run at very low frequency [10, 11], or depend on external high-power computing platforms [13]. For locomotion tasks, the observation-action data are commonly encoded at the timestep level [14, 15, 16, 17], which is straightforward and efficient for producing joint commands in an end-to-end manner. However, it limits the transformer’s direct access to sensory information and still relies on fixed sensor input as they are hidden behind linear projection, thereby constraining its in-context understanding capability and multimodal nature of the data.
@@
To address the above limitations, we propose Masked Sensory-Temporal Attention (MSTA), a novel transformer-based model for end-to-end quadruped locomotion control. It achieves sensor input generalization with its multimodal nature, while still being directly deployable on physical systems. Specifically, in MSTA, all sensory data are discretized and tokenized to form a long proprioceptive information sequence. Inspired by the work [18] on learning spatiotemporal information in video understanding, a random mask is applied to remove a portion of the observation during training. This significantly enhances the model’s sensory-temporal understanding, to better handle different combinations of sensor data and serve as a foundation for incorporating unseen data. Additionally, it aids in identifying the most essential sensory information, thereby reducing the computational power required for physical deployment.
We conduct extensive experiments in the simulation and physical world. Evaluation results demonstrate that MSTA can efficiently handle incomplete sensory information, even with half of the data missing. It is also robust against unseen data, making it a solid foundation for further extensions. With direct sensory-temporal attention, the model is flexible enough to mix-and-match desired information for finetuneing, meeting the requirement for different end-to-end quadruped locomotion control in the physical world.
II Related Work
II-A Sim-to-Real Policy Learning in Legged Locomotion
@@
Reinforcement learning (RL) has gained significant attention in developing robotic controllers for tasks such as legged locomotion [1, 2, 3, 4, 15], eliminating the need for extensive prior knowledge. With the advance of robotic simulation, RL-based locomotion is often trained in virtual environments [4, 19, 20] with diverse terrains [20] and randomized environmental factors [3, 4] to improve the policy robustness. This technique is commonly known as domain randomization (DR). Training in simulators also provides rich information, some of which is not easily accessible in the real world (i.e., privileged information). To better interpret such information and bridge the sim-to-real gap when deploying policies to the physical world, system identification is commonly used to transfer knowledge to a deployable student policy. For instance, Lee et al. [3] employed action imitation to infer teacher behaviors using historical proprioceptive data. Kumar et al. [4] further developed a two-stage adaption framework for faster and more robust online transfer, which has become the foundation for many subsequent works [21, 15]. Another approach combines transfer loss with RL loss for joint optimization [22, 16] to allow the student to explore with teacher guidance to the maximization of reward return.
II-B Transformer in Robotics
@@
The transformer-based models have been introduced to solve robotic tasks. For instance, Decision Transformer [23] converts states, actions, and rewards into embeddings using an encoder. Trajectory Transformer [24] uses the complete discretized trajectory for language model-like autoregressive prediction. Building on these frameworks, Gato [9] was developed to serve as a general agent for hundreds of tasks, including real-world robotic manipulation. Vision Language Models (VLM) [10] and Vision Language Action Models (VLA) [11, 12] use transformers as interfaces for scene and language understanding to provide high-level commands for robotic control and human-robot interaction but due to the enormous model size, they often run at only 5 Hz [10, 11].
@@
For legged locomotion, which requires real-time control, some work involves outputting high-level commands as an interface. For instance, Yang et al. [25] developed a transformer model for vision-based locomotion, which outputs high-level velocity commands and relies on a dedicated low-level controller for motor control. Similarly, Tang et al. [26] uses gait pattern as the interface for a low level controller. The external controller often requires additional design and training, and to bring transformers to direct motor control, Lai et al. [15] proposed TERT, which utilizes historical observation-action pairs to generate target motor commands directly. Barkour [14] uses a similar architecture to merge multiple specialist policies into a single locomotion policy. Radosavovic et al. [16] applied a similar method to the bipedal locomotion task and later reformulated it as a next token prediction problem [17]. Recently, Sferrazza et al. proposed BoT [27], an embodiment-aware transformer network with body-induced bias based on the embodiment graph with embody-specific masking. However, to achieve real time onboard inference, end-to-end controllers relies on fixed information on each timestep and node for ecoding, exhibiting inflexibility of handling different inputs.
III Preliminary
@@
We adopt the two-stage teacher-student transfer approach from TERT [15] as the basis, which utilizes a well-trained teacher policy through RL with privileged information.
III-A Simulation Environment
@@
We implement the simulation environment based on Isaac Gym and its open-source library IsaacGymEnvs [19] to enable massive parallel training.
Terrain and Curriculum. We adopt the terrain curriculum from [20] with five terrain types (smooth slope, rough slope, stairs up, stairs down, discrete obstacle) and difficulty curriculum. The agent progresses and regresses the level based on the episode cumulative tracked linear reward.
Domain Randomization. To enhance the robustness of the policy, DR is used in the simulation following [20, 28]. We sample the commanded longitudinal and lateral velocity from [-1.0, 1.0] m/s, and horizontal angular velocity first calculated based on sampled heading and capped at [-1.0, 1.0] rad/s. Due to the significant computation required for transformer, a system delay is added [28].
@@
Observations and Actions. The privilege observation $e_{t}$ for teacher training contains ground-truth data gathered from simulation, including base linear and angular velocity, orientation, surrounding height map and randomized parameters as described above. For proprioceptive information, we use three commonly seen low-level sensors from quadrupeds. i.e., joint encoders, IMU and foot contact sensors. These sensors can provide five sensory data, including joint position $q\in\mathbb{R}^{12}$, joint velocity $\dot{q}\in\mathbb{R}^{12}$, angular velocity $\omega\in\mathbb{R}^{3}$, gravity vector $g\in\mathbb{R}^{3}$ and binary foot contact $c\in\mathbb{R}^{4}$. Furthermore, the randomly sampled user command target $cmd=[v_{x},v_{y},\omega_{z}]$ and actions from previous step $a_{t-1}\in\mathbb{R}^{12}$ are added, resulting in an observation of $o_{t}\in\mathbb{R}^{49}$ for each step. To gather the temporal information, a list of historical proprioceptive information $[o_{0},o_{1},\cdots,o_{T}]$ from past $T=15$ steps is stored. Thus, full observation is in the $\mathbb{R}^{49\times 15}$ space. Both the teacher and student output the desired joint position $a_{t}$, which is further processed by a PD controller for the output torque $\tau=K_{p}(\hat{q}-q)+K_{d}(\hat{\dot{q}}-\dot{q})$, with base stiffness and damping set to 30 and 0.7 respectively and the target joint velocity $\hat{\dot{q}}$ set to 0.
@@
Reward Function for RL. The reward functions are designed to encourage the agent to follow the commanded velocity. Following [20, 4, 28], we primarily penalize the linear and angular movement along other axes, large joint acceleration and excessive power consumption.
Figure 2: Overview of our MSTA. We gather proprioceptive information from commonly seen low-level sensors for discretization and tokenization. Similar to video understanding, we add additional embedding in three dimensions: sensor type, sensor dim and time. Before being passed to the transformer, a random mask is applied to partially remove the information and a learnable state embedding <S> is used to consolidate the information for action prediction. The target joint position output is passed to the PD controller for direct joint control.
III-B Teacher Policy and Training
@@ IV Methodology
We present MSTA, a novel transformer-based model to generate a generalized understanding of low-level proprioceptive information for quadruped locomotion in complex terrains to handle different sensor set equipped on various robot models or when the sensors are damaged and not available. Unlike previous works [15, 16, 17], where each observation-action pair is processed at the timestep level, we treat each sensor modality individually so that the transformer can learn at the lowest level possible. With this foundational understanding, our model is capable of handling different combinations of sensor inputs, enabling better generalization and flexibility. It can potentially be extended to incorporate high-dimensional sensors for more complex tasks. Fig. 2 shows the overview of MSTA.
@@
Sensory-Action Data Tokenize. To learn in-context information at the lowest sensor level, each modality is encoded individually. Instead of the linear projection used in previous transformer-based locomotion controllers [15, 16, 17], where all sensory observations are merged, individual continuous sensor and control data are mapped to tokens directly. Following previous works [9, 10, 11], we pass the normalized data through an encoder to discretizes the value into 256 bins, which are further mapped into a learnable embedding space with $d=128$ dimensions. Compared to timestep level encoding, in this way, the most information is preserved for in-context understanding by transformer.
@@
Positional Embedding and Sensor Type Embedding. We view the encoded information in a three-dimensional way, sensor type, sensor channel and timestep. To distinguish proprioceptive information from different sources with temporal relations, two additional embeddings are added. The first one is a fixed 2D sin-cos position embedding $e_{p}$ [30] applied on the channel dimension and time axis of each sensor. For instance, $e_{P}^{i,t}$ means the embedding added to the i-th channel at timestep $t$ This allows the model to handle sensors with varying lengths of dimensions and historical time windows directly and be easily extendable. To accommodate the multimodal nature of the sensory data, another learnable embedding $e_{S}$ is add to indicate each sensor type. This enables easy mix and match of information from different sensors without concerns about the order or placeholders. When new sensors are added, a new sensor embedding can be trained and added in. Thus, for the embedding of i-th channel of a sensor at timestep $t$, with original encoded token embedding $e_{t}^{a,i}$, the finial value in the sequence $\mathcal{T}$ is:
$\mathcal{T}_{t}^{a,i}=e_{t}^{a,i}+e_{P}^{i,t}+e_{S}^{a}$ (2)
Random Masking. Inspired by the use of masking in image and video understanding [31, 18] with autoencoders to improve vision understanding, we create a binary mask $M$ based on the target ratio $\alpha$ to randomly mask out portions of the collected sensory, which are directly removed tokens from the original sequence:
$\mathcal{T}_{M}=\{e_{i}\in\mathcal{T}:M_{i}=1\}$ (3)
@@
Since only part of the sensory data are visible to the network, the model is required to infer and reconstruct the missing information from them, thereby enhancing its understanding of the relationships between different sensory inputs. Furthermore, random masking significantly reduces the training time and computational resources required. With sensor level tokens, the input sequence length grows from $T$ to $49T$ for observation, and as the complexity of self-attention is necessarily quadratic in the input length [32], the added overhead is enormous, and masking makes it more feasible to run during massive parallel training.
@@
Transformer Model. We implement a vanilla transformer model to process the generated tokens. The model consists of multihead self-attention blocks with an MLP ratio of 2.0. An additional learnable state embedding <S> is added to the end of the masked sequence $\mathcal{T}_{M}$ to consolidate the processed information [33], which is subsequently projected into the action space with an MLP network $\pi$:
$\begin{gathered}l_{t}=\mathrm{\texttt{MSTA}{}}([\mathcal{T}_{M},\texttt{<S>}]),\;\;\;\;a_{t}=\pi(l_{t})\end{gathered}$ (4)
@@
Teacher-Student Transfer. Following TERT [15], we train MSTA with a two-stage transfer strategy. In the first offline pretraining stage, trajectory is gathered by unrolling the well-trained teacher policy while the student will predict the next actions. This is to ensure that the student can produce reasonable actions during the second online correction stage to overcome the gap of distribution shift by training on its own trajectory. We minimize the loss for action prediction: